# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v2

@@ Introduction
Figure 6: Visual comparison between Interactive-CLIP and original CLIP.
Case Study
As shown in Figure 5, we select three examples to further demonstrate the robustness of InterCLIP-MEP. We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high entropy and incorrect outcomes. However, because the InterCLIP-MEP framework leverages the Memory-Enhanced Predictor (MEP) to utilize historical knowledge of test samples during inference, This approach effectively mitigates the issue for cases where the classification module fails to correctly identify the results by using the historical knowledge of the test samples.
Visualization
To further validate that Interactive-CLIP can more effectively capture the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we visualize the areas of focus during the inference process of the visual model in Figure 6. We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, Interactive-CLIP focuses more accurately on the location of the sea lions compared to the original CLIP. In the second example, which complains about traffic congestion, Interactive-CLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
Conclusion and Future Work
In this paper, we propose a novel framework for robust multi-modal sarcasm detection, termed InterCLIP-MEP. We introduce Interactive CLIP (InterCLIP) as the backbone for extracting sample representations. Unlike the original CLIP, InterCLIP embeds representations from different modalities into the encoder, effectively capturing the intricate interactions between text and image. For inference, we introduce a Memory-Enhanced Predictor (MEP) that leverages historical knowledge of test samples to enhance inference, providing a more robust alternative to directly using the classification module for sarcasm recognition. This approach significantly improves the detection of sarcastic samples. Our experimental results demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark. In future work, we plan to extend our approach to sarcasm recognition tasks across additional modalities, such as speech and video.
Acknowledgments
The research described in the paper received funding from the Anhui Provincial Natural Science Foundation through Grant No. 2308085MF220. Additionally, it was co-supported by the Anhui University Natural Science Foundation, which provided funding through Grant Nos. 2022AH050972 and KJ2021A0516.

### v3

@@ Introduction
Figure 6: Visual comparison between InterCLIP and CLIP.
Case Study and Visualization
As shown in Figure 5, we select three examples to further demonstrate the robustness of InterCLIP-MEP. We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high prediction entropy and incorrect outcomes. However, MEP effectively mitigates the issue for cases where $\mathcal{F}_{c}$ fails to correctly identify the results by using the historical knowledge of the test samples. To further validate that InterCLIP can more effectively capture the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we visualize the areas of focus during the inference process of the visual model in Figure 6. We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, InterCLIP focuses more accurately on the location of the sea lions compared to the original CLIP. In the second example, which complains about traffic congestion, InterCLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
Conclusion
In this paper, we propose a novel framework for robust multi-modal sarcasm detection, termed InterCLIP-MEP. We introduce an Interactive CLIP (InterCLIP) as the backbone for extracting sample representations. Unlike the original CLIP, InterCLIP embeds representations from different modalities into the encoder, effectively capturing the intricate interactions between text and image. For inference, we introduce a Memory-Enhanced Predictor (MEP) that leverages historical knowledge of valuable test samples to enhance inference, providing a more robust and reliable multi-modal sarcasm detection. Our experimental results demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark.